Page tree
Skip to end of metadata
Go to start of metadata

I combed the remaining site todos page, the whiteboard, and my own notes to breakdown what I think needs to be done next at each site.  Intended to be rearranged as things get planned for particular days.

As of 7/26, sw3 and se4 are the only sites still needing soil sensors.

Still to do:

ne1:

  • limit switches for raising and lowering are out of adjustment
  • shoot sonic boom angles

nw1:

  • shoot sonic boom angles

nw2 (40' Rohn)

  • shoot sonic boom angles
  • check tipping bucket (unplugged 7/26)

nw3 (PAM):

  • shoot sonic boom angles
  • check tipping bucket (unplugged 7/19)

nw4:

  • shoot sonic boom angles

ne2:

  • shoot sonic boom angles

ne3:

  • shoot sonic boom angles

ne4:

  • shoot sonic boom angles

sw1:

  • 2m trh needs set screw
  • shoot sonic boom angles

sw3:

  • needs mote and soils
  • shoot sonic boom angles

sw4:

  • diagnose low h2o and co2 values - see if cleaning irga increases the values by 10-20%
  • shoot sonic boom angles

se2:

  • shoot sonic boom angles

se3:

  • shoot sonic boom angles

se4 (PAM):

  • soils
  • replace sock on ec150
  • shoot sonic boom angles

se6:

  • shoot sonic boom angles

sw2 (100' Rohn):

  • shoot sonic boom angles

se5 (40' Rohn):

  • shoot sonic boom angles

Software/displays

  • webplots
  • QC charts
  • cockpit in base (done 7/18)
  • ncharts. Serve from NCAR and ustar?


  • No labels